OpenStudy (anonymous):

What is the solution? 3u+2v=-15 2u+v=-8

OpenStudy (anonymous):

3u+2v=-15 is u=-2v/3 -5 2u+v=-8 is u=-1/2v-4

OpenStudy (anonymous):

which is the solution?